OpenStudy (anonymous):

A voltmeter reads 200.0 V between parallel plates that are 2.0 cm apart. What is the strength of the electric field between the two plates? A. 1.0 × 10^1 N/C B. 1.0 × 10^2 N/C C. 1.0 × 10^3 N/C D. 1.0 × 10^4 N/C

OpenStudy (michele_laino):

this realtionship holds: \[\Delta V = Ed\]

OpenStudy (anonymous):

200*2.0 ?

OpenStudy (anonymous):

:/

OpenStudy (michele_laino):

no, we have: \[E = \frac{{\Delta V}}{d} = \frac{{200}}{{2 \times {{10}^{ - 2}}}}\]

OpenStudy (anonymous):

ohh oops!! so we get 10,000 ? so our solution is choice D?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! thanks!!:D

OpenStudy (michele_laino):

:)